An economist estimated that the cost function of a single-product firm is:


C(Q) = 80 + 20Q + 25Q2 + 5Q3.


Based on this information, determine the following:


a. The fixed cost of producing 10 units of output.



b. The variable cost of producing 10 units of output.



c. The total cost of producing 10 units of output.



d. The average fixed cost of producing 10 units of output.



e. The average variable cost of producing 10 units of output.



f. The average total cost of producing 10 units of output.



g. The marginal cost when Q = 10.





1
Expert's answer
2021-11-23T17:21:38-0500

a) The fixed cost of producing 10 units of output.

FC = "\$ 80"

b. The variable cost of producing 10 units of output.

"VC(10) = 20\times 10 + 25\times 10\\^2 + 5\times 10\\^3 =\$ 7700"

Variable cost = "\$ 7700"

c. The total cost of producing 10 units of output.

Total Cost "= FC + VC"

TC = "\$80 +\$ 7700 = \$7780"

d. The average fixed cost of producing 10 units of output.

"AFC(10) = \frac{80}{10} = \$ 8"

e. The average variable cost of producing 10 units of output.

"AVC(10) = \frac{7700}{10} = \$770"

f. The average total cost of producing 10 units of output.

"ATC = 770 + 8 = \$ 778"

g. The marginal cost when Q = 10.

"MC(10) = 20 + (2\times 25)(10) + (3\times5)(10)\\^2= \$2,020"